A family goes on a vacation to a water park. During the drive to the water park the weather caused the family to average 36 miles per hour. After their vacation the weather improved, so they could average 80 miles per hour on the return trip. If the distance each way was equal , and the entire time driving was 2.9 hours, find the distance one way to the water park.
Answer by Cromlix(4381)

Trip to water park averaged 36mph
= D/36
Trip from water park averaged 80mph
= D/80
So, D/36 + D/80 = 2.9
Common denominator = 2880
(80D + 36D)/2880 = 2.9
Cross multiply
116D = 2880 * 2.9
116D = 8352
D = 8352/116
D = 72 miles
Distance one way to water park = 72 miles.
Proof:
D/36 + D/80 = 2.9
72/36 + 72/80 = 2.9
2.9 = 2.9
Hope this helps.
:-)